8. You are the treasurer for your companys actuarial club. The clubs account had a fund balance of $1,050 on January 1 and
Answer #1

Using the Modified dietz approximation method to calculate the dollar weighted rate of return:

Rate of return = (EMV - BMV - c) (no.ofdays fromcash flowtillenddate/Totalnumberofdays) * C;) (BMV +

where

EMV = Ending Market Value = 1160

BMV = Beginning Market Value = 1050

Ci = cash flow after Di days i.e. after 31 days if the first cash flow is on Jan 31.

No. of days from cash flow till end date = 365 - Di = 365 - 31 = 334 for the first cash flow

Total number of days = 365 (since we are considering a year)

The values are calculated in excel and the table is copied below. The formula is copied in the table below:

The answer is option B: 3.23%

Date No. of days Fund value Deposit Withdrawal Mod Dietz Cash flow
01-01-07 364.00 1050
01-31-07 334.00 90 82.35616 90
02-28-07 306.00 90 75.45205 90
03-31-07 275.00 90 67.80822 90
04-30-07 245.00 90 60.41096 90
05-01-07 244.00 600 -401.096 -600
05-31-07 214.00 90 52.76712 90
06-15-07 199.00 400 -218.082 -400
06-30-07 184.00 90 45.36986 90
07-31-07 153.00 90 37.72603 90
08-31-07 122.00 90 30.08219 90
09-30-07 92.00 90 22.68493 90
10-31-07 61.00 90 15.0411 90
11-30-07 31.00 90 7.643836 90
12-31-07 0.00 1160 90 0 90
Total -121.836 80
Return 3.232%
Date No. of days Fund value Deposit Withdrawal Mod Dietz Cash flow
39083 =DAYS($A$39,A3) 1050
39113 =DAYS($A$39,A4) 90 =(D4-E4)*(B4/365) =D4-E4
39141 =DAYS($A$39,A5) 90 =(D5-E5)*(B5/365) =D5-E5
39172 =DAYS($A$39,A6) 90 =(D6-E6)*(B6/365) =D6-E6
39202 =DAYS($A$39,A7) 90 =(D7-E7)*(B7/365) =D7-E7
39203 =DAYS($A$39,A8) 600 =(D8-E8)*(B8/365) =D8-E8
39233 =DAYS($A$39,A9) 90 =(D9-E9)*(B9/365) =D9-E9
39248 =DAYS($A$39,A10) 400 =(D10-E10)*(B10/365) =D10-E10
39263 =DAYS($A$39,A11) 90 =(D11-E11)*(B11/365) =D11-E11
39294 =DAYS($A$39,A12) 90 =(D12-E12)*(B12/365) =D12-E12
39325 =DAYS($A$39,A13) 90 =(D13-E13)*(B13/365) =D13-E13
39355 =DAYS($A$39,A14) 90 =(D14-E14)*(B14/365) =D14-E14
39386 =DAYS($A$39,A15) 90 =(D15-E15)*(B15/365) =D15-E15
39416 =DAYS($A$39,A16) 90 =(D16-E16)*(B16/365) =D16-E16
39447 =DAYS($A$39,A17) 1160 90 =(D17-E17)*(B17/365) =D17-E17
Total =SUM(F4:F17) =SUM(G3:G17)
Return =(C17-C3-G18)/(C3+F18)